The sum of the digits of a two-digit number is 8. The result of subtracting the units digit from the tens digit is -4. Define the variables and write the system of equations that can be used to find the number. Then solve the system and find the number.


1
Expert's answer
2021-01-05T16:28:06-0500

Solution. Represented two-digit number as xy. According to the condition of the problem


"x+y=8"

also


"x-y=-4"

As a result, we obtain the system of equations

We solve the system of equations. Let us add the first and second equations of the system


"x+y+x-y=8-4"

Hence, get


"2x=4"

"x=2"

Substituting the obtained value of x into the first equation, we obtain


"2+y=8""y=6"

As a result, the number is


"26"

Answer: 26.